# Empowering Targeted Neighborhood Search via Hyper Tour for Large-Scale TSP

### v2

@@ 4.1 Experimental Settings
Datasets. We first evaluate our approach on the standard TSP benchmarks (fu2021generalize; luoboosting; qiu2022dimes; li2025destroy; zhou2025dualopt), uniformly generated within a unit square (128 instances with NN=1K; 32 with NN=5K; 16 each with NN=10K, 20K, 50K). Beyond standard benchmarks, we further evaluate the generalization of HyperNS on real-world and cross-distribution datasets. We use TSPLIB (reinelt1991tsplib) with diverse instances, including original TSPLIB (21 instances, 150–18,512 cities), Nationallib (18 instances, 194–71,009 cities), and VLSI (84 instances, 131–52,057 cities)—as well as cross-distribution datasets from INVIT (fang2024invit) with four node distributions (uniform, clustered, explosion, implosion), enabling a thorough evaluation of robustness and generalization.
@@
Table 2: Performance on TSPLIB (reinelt1991tsplib).
  Size LKH3 (helsgaun2017extension) DRHG (li2025destroy) DualOpt (zhou2025dualopt) HyperNS
@@
Comparison methods. We compare our algorithm against a range of TSP solvers: (1) classic solvers like Concorde (cook2011traveling) and LKH3 (helsgaun2017extension) [^4] LKH3 is a highly specialized, manually crafted solver that incorporates extensive domain knowledge and sophisticated heuristics, such as k-opt, $\alpha$-measure, partitioning and tour merging methods, iterative partial transcription, and backbone-guided search., and (2) neural-based methods such as AM-greedy (kool2018attention), GCN-BS (joshi2019efficient), AttGCRN+MCTS (fu2021generalize), DIMES (qiu2022dimes), DIFUSCO (sun2023difusco), SoftDist (xia2024position), Select&Optimize (SO) (cheng2023select), UDC (zhengudc), DRHG (li2025destroy), INViT (fang2024invit), PRC (luoboosting), and DualOpt (zhou2025dualopt). All baselines are run with default settings, except Select&Optimize, which we reimplemented due to the unavailable source code. Since our main focus is on integrating neural and traditional methods, we primarily compare our HyperNS with similar approaches and list the results of LKH3 for indicative purposes following previous neural methods.
Implementation details. Our experiments were conducted on an Intel Xeon Gold 6148 CPU@2.40GHz and an NVIDIA Tesla V100 PCIe 32GB GPU. The GPU was used for neural models, and the CPU for conventional solvers. Att-GCRN-GI [^5] The trained model is fixed and directly applied to TSP instances of varying sizes and distributions without further retraining. is trained on 900K instances with nn=100100, following the standard configuration in prior works (joshi2019efficient; fu2021generalize). The subgraph size $p$ and the number of vertices per grid $\gamma$ for sparse heatmap graph generation are set to 100 and 30, respectively, to match Att-GCRN-GI’s training graph sizes. $\alpha$ for targeted neighborhood search is set to 1000 to maintain consistent magnitudes across different terms of the score. Based on the experimental results in Exp-6, we set the nearest neighbor number $m$ to 100, sub-tour length $l_{s}$ to 100 and the number of iterations $I_{3}$ for sub-tour optimization to 2.
Table 3: Results on large-scale real-world instances (reinelt1991tsplib). For neural methods, only DualOpt is reported, as it achieves the best results among all neural baselines (Tables 1 & 2).
@@
Table 4: Performance on non-uniform datasets (fang2024invit).
@@
DRHG (li2025destroy) 0.31 2.31h 0.88 3.13h 1.33 4.26h 0.49 2.33h 1.35 3.16h 2.14 4.27h
INViT (fang2024invit) 5.99 0.26h 6.46 0.76h 6.06 0.73h 8.63 0.26h 8.57 0.76h 8.79 0.73h
DualOpt (zhou2025dualopt) 0.82 0.31h 1.3 0.64h 1.18 0.87h 1.07 0.31h 1.68 0.67h 2.32 0.93h
@@
DRHG (li2025destroy) 0.43 2.31h 1.47 3.14h 2.53 4.29h 0.41 2.31h 1.32 3.13h 2.48 4.26h
INViT (fang2024invit) 8.57 0.26h 9.43 0.76h 9.05 0.74h 6.35 0.26h 7.41 0.76h 6.21 0.73h
DualOpt (zhou2025dualopt) 1.01 0.32h 1.54 0.65h 1.86 0.91h 0.84 0.33h 1.47 0.66h 1.75 0.88h
@@ 4.2 Experimental Results
Exp-4: Performance on Non-Uniform Instances. We evaluate the generalization ability of our method on cross-distribution instances with DRHG (li2025destroy), INViT (fang2024invit) and DualOpt (zhou2025dualopt). The results are shown in Table 4. Our HyperNS achieves the lowest gap among these methods on all datasets, showing its great cross-distribution generalizability. Specifically, the average gaps across different distributions produced by HyperNS are 0.59%, 0.86%, 0.86%, and 0.69% lower than those of DualOpt. Additionally, the relative gap increase from uniform distribution to other distributions for HyperNS on TSP10K is much lower than DualOpt (38.0% v.s. 68.5% on average). This indicates that our proposed method possesses both excellent cross-distribution and cross-size generalizability.
@@
Exp-5: Ablation Study. We further investigate the impact of different components in our model, including subgraph sampling, hyper tour, and hyper tour-guided initialization and optimization. The results are shown in Table 5. First, we test the effectiveness of Att-GCRN-GI and replace it with Att-GCRN, shown as ID-1. Our Att-GCRN-GI outperforms Att-GCRN, which brings an average 0.10% decrease in the average gap. Second, we test the effectiveness of “edge covering” for subgraph sampling, and replace it with the “node covering” from (fu2021generalize), shown as ID-2. Our edge covering method outperforms the node covering method, which brings an average 0.25% decrease in the average gap and 4.84% decrease in running time on all datasets. Next, we compare the effects of our hyper tour-guided initialization (HTI) with different initialization methods, including the traditional “greedy” and “beam search (BS)” methods. We also test the results of randomly generating the order of supernodes without hyper tour guidance, denoted as “w/o. HT”. The results are shown as IDs-3 to 5. Compared with the widely adopted greedy initialization method, our hyper tour-guided initialization reduces the average gap by 88.9% , with total running times increasing from 3.21 to 4.06 hours. This increase in running time is acceptable given the significant reduction in the average gap. In contrast, the beam search-based initialization is substantially slower and could not be applied to TSP10K and TSP20K instances. Besides, the introduction of the hyper tour on our tour initialization reduces the average gap by 69.44% and running time by 18.31% across all datasets, validating its effectiveness in improving convergence and overall results. These findings justify the design of our hyper tour-guided initialization method, which improves convergence speed and leads to better overall results.
@@ A.1 Details of Related Work
Traditional TSP solvers typically fall into three categories: exact methods (cook2011traveling), approximation methods (dumitrescu2003approximation), and heuristic methods (helsgaun2000effective). With the rise of neural networks, recent research increasingly explores neural-based TSP methods or approaches that combine traditional solvers with neural networks. Early neural methods, typically limited to instances with up to 1,000 vertices, can be divided into constructive-based methods (vinyals2015pointer; bello2016neural; deudon2018learning; kool2018attention; khalil2017learning; li2024fast; grinsztajn2023winner; chalumeau2024combinatorial; luo2024neural; kwon2020pomo; jin2023pointerformer; drakulic2023bq; lyu2024scalable; jacobs21reinforcement; zhang2022learning; jiang2022learning; bi2022learning; zhou2023towards; son2023meta; li2025unify; wang2024asp) that generate tours from scratch with neural models and improvement-based methods (joshi2019efficient; xiao2024distilling; li2023distribution; min2024unsupervised; chen2019learning; wu2021learning; lu2019learning; hudson2022graph; yao22reversible; ma2023learning; ye2023deepaco; wang2025efficient; li2025unify) that first generated a tour and then iteratively refined under neural guidance toward better solutions.
@@
Constructive-based Methods. These methods generate TSP tours from scratch using neural models. The Pointer Network (vinyals2015pointer) first introduced an encoder-decoder framework for supervised incremental solution generation. Later advancements incorporated reinforcement learning (bello2016neural; deudon2018learning; kool2018attention; khalil2017learning), diffusion models (li2024fast), refined decoders (grinsztajn2023winner; chalumeau2024combinatorial; luo2024neural), utilized TSP tour symmetry (kwon2020pomo; jin2023pointerformer; drakulic2023bq) and the nearest neighbors (lyu2024scalable), trained on diverse instance distributions (jacobs21reinforcement; zhang2022learning; jiang2022learning; bi2022learning), explored meta-learning (zhou2023towards; son2023meta), joint probability estimation (li2025unify) and curriculum learning (wang2024asp). However, these approaches are limited to TSP instances with up to 1,000 vertices and struggle with larger instances.
@@
Improvement-based Methods. These methods combine neural models with traditional search methods, and can be classified into two types. The first type is to generate the heatmap matrix (probability of an edge being in the optimal solution) based on the GNN model and then use search methods like beam search (joshi2019efficient; xiao2024distilling), 2-opt (li2023distribution) and best-first local search (min2024unsupervised) to optimize the solution. Another type of search-based methods adopts neural models (mainly reinforcement learning) to guide the local search, including 2-opt (chen2019learning; wu2021learning; lu2019learning; hudson2022graph; yao22reversible), k-opt (ma2023learning), beam search (choo2022simulation) and Ant Colony Optimization (ye2023deepaco). The lack of optimal labels for large-scale instances makes training neural network models challenging, which restricts these methods to handling instances with up to 1,000 vertices.
@@ A.2.1 Pretrained Encoder-Decoder Framework for the Geometric Information
Inspired by the work of (zhou2023improving), we construct a pretrained encoder-decoder network and incorporate the geometric information-aware hidden representations from the encoder as part of the node features for Att-GCRN. Specifically, our model uses an encoder that comprises an MLP layer to map the coordinates of vertices in $\mathbb{V}$ into a high-dimensional vector space, followed by layer normalization. This process can be formalized as follows:
@@ A.3 Time Complexity Analysis
Our method comprises five key components: sparse heatmap graph construction, hyper tour generation, initialization, targeted neighborhood search, and sub-tour optimization, with GCRN model and LK search employed to address various subproblems. The complexity of GCRN model is $O(n^{2})$ (Wu21Comprehensive), while LK search is a PLS-complete problem (papadimitriou1992complexity), and its complexity cannot be precisely defined but is denoted as $O(\texttt{L}(n))$.

### v3

@@ 4.1 Experimental Settings
Datasets. We first evaluate our approach on the standard TSP benchmarks (Fu et al., 2021; Luo et al., 2025; Qiu et al., 2022; Li et al., 2025a; Zhou et al., 2025), uniformly generated within a unit square (128 instances with NN=1K; 32 with NN=5K; 16 each with NN=10K, 20K, 50K). Beyond standard benchmarks, we further evaluate the generalization of HyperNS on real-world and cross-distribution datasets. We use TSPLIB (Reinelt, 1991) with diverse instances, including original TSPLIB (21 instances, 150–18,512 cities), Nationallib (18 instances, 194–71,009 cities), and VLSI (84 instances, 131–52,057 cities)—as well as cross-distribution datasets from INVIT (Fang et al., 2024) with four node distributions (uniform, clustered, explosion, implosion), enabling a thorough evaluation of robustness and generalization.
@@
Table 2: Performance on TSPLIB (Reinelt, 1991).
  Size LKH3 (Helsgaun, 2017) DRHG (Li et al., 2025a) DualOpt (Zhou et al., 2025) HyperNS
@@
Comparison methods. We compare our algorithm against a range of TSP solvers: (1) classic solvers like Concorde (Cook et al., 2011) and LKH3 (Helsgaun, 2017) [^4] LKH3 is a highly specialized, manually crafted solver that incorporates extensive domain knowledge and sophisticated heuristics, such as k-opt, $\alpha$-measure, partitioning and tour merging methods, iterative partial transcription, and backbone-guided search., and (2) neural-based methods such as AM-greedy (Kool et al., 2019), GCN-BS (Joshi et al., 2019), AttGCRN+MCTS (Fu et al., 2021), DIMES (Qiu et al., 2022), DIFUSCO (Sun and Yang, 2023), SoftDist (Xia et al., 2024), Select&Optimize (SO) (Cheng et al., 2023), UDC (Zheng et al., 2024), DRHG (Li et al., 2025a), INViT (Fang et al., 2024), PRC (Luo et al., 2025), and DualOpt (Zhou et al., 2025). All baselines are run with default settings, except Select&Optimize, which we reimplemented due to the unavailable source code. Since our main focus is on integrating neural and traditional methods, we primarily compare our HyperNS with similar approaches and list the results of LKH3 for indicative purposes following previous neural methods.
Implementation details. Our experiments were conducted on an Intel Xeon Gold 6148 CPU@2.40GHz and an NVIDIA Tesla V100 PCIe 32GB GPU. The GPU was used for neural models, and the CPU for conventional solvers. Att-GCRN-GI [^5] The trained model is fixed and directly applied to TSP instances of varying sizes and distributions without further retraining. is trained on 900K instances with nn=$100$, following the standard configuration in prior works (Joshi et al., 2019; Fu et al., 2021). The subgraph size $p$ and the number of vertices per grid $\gamma$ for sparse heatmap graph generation are set to 100 and 30, respectively, to match Att-GCRN-GI’s training graph sizes. $\alpha$ for targeted neighborhood search is set to 1000 to maintain consistent magnitudes across different terms of the score. Based on the experimental results in Exp-6, we set the nearest neighbor number $m$ to 100, sub-tour length $l_{s}$ to 100 and the number of iterations $I_{3}$ for sub-tour optimization to 2.
Table 3: Results on large-scale real-world instances (Reinelt, 1991). For neural methods, only DualOpt is reported, as it achieves the best results among all neural baselines (Tables 1 & 2).
@@
Table 4: Performance on non-uniform datasets (Fang et al., 2024).
@@
DRHG (Li et al., 2025a) 0.31 2.31h 0.88 3.13h 1.33 4.26h 0.49 2.33h 1.35 3.16h 2.14 4.27h
INViT (Fang et al., 2024) 5.99 0.26h 6.46 0.76h 6.06 0.73h 8.63 0.26h 8.57 0.76h 8.79 0.73h
DualOpt (Zhou et al., 2025) 0.82 0.31h 1.3 0.64h 1.18 0.87h 1.07 0.31h 1.68 0.67h 2.32 0.93h
@@
DRHG (Li et al., 2025a) 0.43 2.31h 1.47 3.14h 2.53 4.29h 0.41 2.31h 1.32 3.13h 2.48 4.26h
INViT (Fang et al., 2024) 8.57 0.26h 9.43 0.76h 9.05 0.74h 6.35 0.26h 7.41 0.76h 6.21 0.73h
DualOpt (Zhou et al., 2025) 1.01 0.32h 1.54 0.65h 1.86 0.91h 0.84 0.33h 1.47 0.66h 1.75 0.88h
@@ 4.2 Experimental Results
Exp-4: Performance on Non-Uniform Instances. We evaluate the generalization ability of our method on cross-distribution instances with DRHG (Li et al., 2025a), INViT (Fang et al., 2024) and DualOpt (Zhou et al., 2025). The results are shown in Table 4. Our HyperNS achieves the lowest gap among these methods on all datasets, showing its great cross-distribution generalizability. Specifically, the average gaps across different distributions produced by HyperNS are 0.59%, 0.86%, 0.86%, and 0.69% lower than those of DualOpt. Additionally, the relative gap increase from uniform distribution to other distributions for HyperNS on TSP10K is much lower than DualOpt (38.0% v.s. 68.5% on average). This indicates that our proposed method possesses both excellent cross-distribution and cross-size generalizability.
@@
Exp-5: Ablation Study. We further investigate the impact of different components in our model, including subgraph sampling, hyper tour, and hyper tour-guided initialization and optimization. The results are shown in Table 5. First, we test the effectiveness of Att-GCRN-GI and replace it with Att-GCRN, shown as ID-1. Our Att-GCRN-GI outperforms Att-GCRN, which brings an average 0.10% decrease in the average gap. Second, we test the effectiveness of “edge covering” for subgraph sampling, and replace it with the “node covering” from (Fu et al., 2021), shown as ID-2. Our edge covering method outperforms the node covering method, which brings an average 0.25% decrease in the average gap and 4.84% decrease in running time on all datasets. Next, we compare the effects of our hyper tour-guided initialization (HTI) with different initialization methods, including the traditional “greedy” and “beam search (BS)” methods. We also test the results of randomly generating the order of supernodes without hyper tour guidance, denoted as “w/o. HT”. The results are shown as IDs-3 to 5. Compared with the widely adopted greedy initialization method, our hyper tour-guided initialization reduces the average gap by 88.9% , with total running times increasing from 3.21 to 4.06 hours. This increase in running time is acceptable given the significant reduction in the average gap. In contrast, the beam search-based initialization is substantially slower and could not be applied to TSP10K and TSP20K instances. Besides, the introduction of the hyper tour on our tour initialization reduces the average gap by 69.44% and running time by 18.31% across all datasets, validating its effectiveness in improving convergence and overall results. These findings justify the design of our hyper tour-guided initialization method, which improves convergence speed and leads to better overall results.
@@ A.1 Details of Related Work
Traditional TSP solvers typically fall into three categories: exact methods (Cook et al., 2011), approximation methods (Dumitrescu and Mitchell, 2003), and heuristic methods (Helsgaun, 2000). With the rise of neural networks, recent research increasingly explores neural-based TSP methods or approaches that combine traditional solvers with neural networks. Early neural methods, typically limited to instances with up to 1,000 vertices, can be divided into constructive-based methods (Vinyals et al., 2015; Bello et al., 2017; Deudon et al., 2018; Kool et al., 2019; Khalil et al., 2017; Li et al., 2024; Grinsztajn et al., 2023; Chalumeau et al., 2023; Luo et al., 2024; Kwon et al., 2020; Jin et al., 2023; Drakulic et al., 2023; Lyu et al., 2024; Jacobs et al., 2021; Zhang et al., 2022; Jiang et al., 2022; Bi et al., 2022; Zhou et al., 2023a; Son et al., 2023; Li et al., 2025b; Wang et al., 2024) that generate tours from scratch with neural models and improvement-based methods (Joshi et al., 2019; Xiao et al., 2024; Li et al., 2023; Min et al., 2024; Chen and Tian, 2019; Wu et al., 2021a; Lu et al., 2019; Hudson et al., 2022; Yao et al., 2022; Ma et al., 2023; Ye et al., 2023; Wang et al., 2025; Li et al., 2025b) that first generated a tour and then iteratively refined under neural guidance toward better solutions.
@@
Constructive-based Methods. These methods generate TSP tours from scratch using neural models. The Pointer Network (Vinyals et al., 2015) first introduced an encoder-decoder framework for supervised incremental solution generation. Later advancements incorporated reinforcement learning (Bello et al., 2017; Deudon et al., 2018; Kool et al., 2019; Khalil et al., 2017), diffusion models (Li et al., 2024), refined decoders (Grinsztajn et al., 2023; Chalumeau et al., 2023; Luo et al., 2024), utilized TSP tour symmetry (Kwon et al., 2020; Jin et al., 2023; Drakulic et al., 2023) and the nearest neighbors (Lyu et al., 2024), trained on diverse instance distributions (Jacobs et al., 2021; Zhang et al., 2022; Jiang et al., 2022; Bi et al., 2022), explored meta-learning (Zhou et al., 2023a; Son et al., 2023), joint probability estimation (Li et al., 2025b) and curriculum learning (Wang et al., 2024). However, these approaches are limited to TSP instances with up to 1,000 vertices and struggle with larger instances.
@@
Improvement-based Methods. These methods combine neural models with traditional search methods, and can be classified into two types. The first type is to generate the heatmap matrix (probability of an edge being in the optimal solution) based on the GNN model and then use search methods like beam search (Joshi et al., 2019; Xiao et al., 2024), 2-opt (Li et al., 2023) and best-first local search (Min et al., 2024) to optimize the solution. Another type of search-based methods adopts neural models (mainly reinforcement learning) to guide the local search, including 2-opt (Chen and Tian, 2019; Wu et al., 2021a; Lu et al., 2019; Hudson et al., 2022; Yao et al., 2022), k-opt (Ma et al., 2023), beam search (Choo et al., 2022) and Ant Colony Optimization (Ye et al., 2023). The lack of optimal labels for large-scale instances makes training neural network models challenging, which restricts these methods to handling instances with up to 1,000 vertices.
@@ A.2.1 Pretrained Encoder-Decoder Framework for the Geometric Information
Inspired by the work of (Zhou et al., 2023b), we construct a pretrained encoder-decoder network and incorporate the geometric information-aware hidden representations from the encoder as part of the node features for Att-GCRN. Specifically, our model uses an encoder that comprises an MLP layer to map the coordinates of vertices in $\mathbb{V}$ into a high-dimensional vector space, followed by layer normalization. This process can be formalized as follows:
@@ A.3 Time Complexity Analysis
Our method comprises five key components: sparse heatmap graph construction, hyper tour generation, initialization, targeted neighborhood search, and sub-tour optimization, with GCRN model and LK search employed to address various subproblems. The complexity of GCRN model is $O(n^{2})$ (Wu et al., 2021b), while LK search is a PLS-complete problem (Papadimitriou, 1992), and its complexity cannot be precisely defined but is denoted as $O(\texttt{L}(n))$.